How People Use Symbols

Marc Horowitz 21.512 May 11, 1992

Human beings have traditionally used symbols of all kinds as a vehicle for understanding and controlling the world around them. Symbols can apply to individuals, groups, or society. A symbol can often make a point to a large number of people in a society, where a factual description might not be convincing enough.

One example of symbols in society is the tropes of everyday language. Within a society, a trope can carry a great deal of direct and subtle meaning. I may leave a test and say to a friend, “That test was a piece of cake!” This means more than merely saying “That test was easy!” The test was so easy I enjoyed taking it, as I might enjoy eating a piece of cake. Tropes in language can also be used to separate a group of people. Almost any subculture has its own set of metaphors which it uses to describe events within that subculture. Someone who understands this language is an insider; anyone else is an outsider. Sporting enthusiasts, political organizations, scientists, doctors, and lawyers are only a few examples of subcultures which have their own symbols in language.

Symbols also play an active role in the day-to-day lives of people. In classical Greek mythology, certain gods, places, or characters are symbolic of certain traits: Zeus is symbolic of power, Aphrodite of beauty, Athena of wisdom, etc. An individual might invoke one of these gods if he or she wanted intervention from that god on his or her behalf. A ruler might sacrifice something to Zeus in hope that he would remain in power. A scholar might keep around something symbolic of Athena in hopes that he might grow wiser, and to indicate his profession to others. Metaphors serve a similar purpose in our culture. Flowers can be a sign of many things. We send them to sick people as a symbol of the health and vitality we wish them to regain. We send them as tokens of friendship or affection. Sometimes this symbolism is extremely ritualized, such as a red rose for love, or a white rose for friendship. Symbols such as flowers or religious icons represent a thought which would be difficult to express otherwise.

Another common symbol is physical appearance. An individual’s appearance can be symbolic for societal rank, occupation, etc. We can see examples of this in Raymond Firth’s essay “Hair as Private Asset and Public Symbol”. In our society, we see hair as an indicator of personality. A male who views himself as a liberal or a non-conformist will grow his hair long. A woman with unshaven legs will be usually be labelled as a radical feminist or something similar. In other societies, the exact form of the symbol may be different, but it will be there. “What is particularly important is that these styles are often not just aesthetic modulations or a matter of individual preference; they are strongly conventionalized, given moral approval and used as instruments of social expression and social control.”^{1} A Native American woman will wear her hair differently if she is married or unmarried. All of these are examples of how a person may alter his or her physical appearance symbolically. The body itself is not the only vehicle for symbolic displays of appearance. A monarch’s bejewelled crown and sceptre are symbolic of his power; nobody but a powerful ruler could afford to waste gold or jewels on a headpiece or a stick. Some symbols are more arbitrary. A bride wears white, metaphorical of an ideal of purity. A mourner wears black, metaphorical of death. In all of these examples, the purpose of the symbol is to set aside or distinguish an individual in a well-understood way.

Symbols do not always play such a passive role. Rituals almost always have a strong symbolic component. Often, an everyday item or place will play a special role in a ritual, symbolizing by its use or form something much more profound. Victor Turner gives several detailed examples in “Ritual Symbolism, Morality, and Social Structure among the Ndembu”. He gives as examples two kinds of trees which are, presumably, common. One kind of tree is the mukala tree, which secretes a red gum. This red gum is symbolic of blood in Ndembu rituals, including, for example, the male circumcision ritual. In this case, the tree is used in the ritual after the circumcision has taken place. The intent is that the wounds should heal quickly, as the gum hardens quickly as well. The purpose of this sort of ritual is to draw a connection between human society and natural order.

1Firth, p. 271

Another common use of symbols is to provide a parallel in nature of something arbitrary in society. We can see this in animal metaphors. The phrase “king of the jungle” demonstrates the lion as a symbol of dominance and power. This well-known symbol might cause a monarch to choose a name such as “Richard the Lion-Hearted”, which would be implicitly understood. In “fox hunting as ritual”, by James Howe, we see examples of animal symbolism. The hounds of the hunter symbolize the hunter, the hunted animal symbolizes the enemy of the hunter, and the hunt itself might symbolize a battle. An English gentleman would treat his hounds and horses with extreme care, breeding them and keeping them as he might his own family or himself. A well-kept hunting pack was an outward symbol of a rich and important hunter. The animal being hunted was a symbolic enemy. Howe mentions “the concept of the noble adversary whose sterling qualities confirm one’s own.”^{2} The most noble animals, which may have been the wild boor, deer, or fox, depending on the time, symbolized the most honorable chase.

In a few weeks, MIT will be holding a graduation ritual in which many symbols can be found. The diploma is a symbol of the knowledge which the student has acquired. Clothing is symbolic of rank. Graduates wear a simple black gown. Higher degrees come with more elaborate colored hoods. (Even the colors are symbolic of the kind of degree.) Professors, deans, and the President will wear special regalia setting them aside from all the other people present. The symbols here are intended to make concrete orderings and distinctions which may seem arbitrary to someone unfamiliar with acedemia. One can look at a B.S. and a Ph.D., and it is clear who has the more significant ranking in the academic heirarchy. The entire ceremony itself is a symbol of passage. Before the ceremony, the person is a student, not yet fully learned in his field. After, he is a scholar, fully knowledgeable in his field, certified to use his skills in the real world, or to go on to more advanced fields of learning.

2Howe, p. 293

People create symbols to add structure to parts of their world which are poorly understood, or to make difficult concepts easier to understand. The symbols of a society provide a common experience base for people to draw from both in everyday life, and in ritual situations. The symbols which can provide a cohesiveness within a society may also serve the function of differentiating outsiders, forcing them to conform to local standards, or be shunned. Symbols add an extra dimension to human life which would be difficult to achieve in any other way. These symbols are intrinsic to, and in many cases, defining of a culture.